OpenStudy (anonymous):

what is the slope of a line perpendicular to rhe line whose equation is y=2x+5?

OpenStudy (anonymous):

the slope of \[\huge y=\color{red}mx+b\] is \(\large \color{red}m\) the slope of the perpendicular line is \[\huge -\frac{1}{\color{red}m}\]

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

jIs the data set “names of the players on the golf team” quantitative or qualitative? If it is quantitative, is it discrete or continuous?

OpenStudy (anonymous):

m1*m2=-1 m1= slope of 1st line m2=slope of the line perpendicular to the 1st line. here eqn is 2x-y+5=0 so m1= 2 therefore m2= -1/2 hope it will help you.